OpenStudy (justin01):

Solve by factoring or finding square roots. 2x^2-3x=0

jhonyy9 (jhonyy9):

like a first step facoriz out the x - what will get ?

jhonyy9 (jhonyy9):

factorize out - sorry

OpenStudy (will.h):

Do you know how to use the GCF. Great common factor?

OpenStudy (pythagoras123):

\[2x^2-3x=x(2x-3)=0\] By zero factor property, x = 0 OR 2x - 3 = 0 x = 0 OR x = 1.5